### v1

@@ IV-D Computation time
1.8/6.9 1.6/1.8 2.9/3.3 4.3/5.1 2.1/8.2 1.9/2.1 3.4/3.8 5.1/5.8
TABLE III: Query/Insert Time ($\mu$s)
IV-E Security analysis
Here, we establish the security of Neonpool, focusing on two main aspects: 1. whether errors made by Neonpool compromise its security or that of the broader network; and 2. Neonpool’s resilience to adversarial attacks. Neonpool uses bloom filter variants which yield false positives and false negatives. However, their effect on transaction propagation vanishes at the network level. This is because each node initializes its filters using a random 128-bit salt, making filters at each node independent and the false positives at each node different. For instance, 99.99% accuracy or an FPR of $0.0001$ at each Neonpool node, implies that the probability that two nodes drop the same transaction is $(0.0001)^{2}$, which is negligible. The same argument applies to false negatives.
@@ IV-E Security analysis
Additionally, an adversary may: 1. trigger false positives to censor specific transactions; 2. craft invalid transactions that evade verification and validation; 3. generate dust or spam.
@@
Literature shows that any bloom filter can be efficiently transformed to be adversarial resilient by applying a pseudo-random permutation of the input [14] i.e. applying a sufficiently large (128-bit) random salt before forwarding it to the bloom filter. This change requires little overhead and randomizes the adversary’s queries by applying a pseudorandom permutation to them; then we may consider the transactions sent by the attacker as random and not as chosen adaptively by the adversary. It is also recommended that a node regenerate its 128-bit random salt every time a new bloomtxFilter or dstxFilter is generated. The adversary has only oracle access to the bloom filter and does not know its internal memory representation or seed.
@@
Secondly, Eve may craft invalid transactions that evade verification and validation i.e. attempt to double-spend. Neonpool preserves the transaction verification and validation mechanism of Bitcoin and Ethereum. With regards to transactions that have conflicting inputs or out-of-order nonce, typically nodes accept and forward the first seen transaction and the first seen can differ for nodes. It is the job of miners to not add conflicting transactions to a block and the network nodes to screen incoming blocks. Since Neonpool maintains complete UTXO and Trie information, Neonpool nodes will reject blocks that include double-spend transactions.
@@
Thirdly, Eve might launch a dust and spam attack or simply replay transactions. Dust and spam frequently occur in the transaction. A 2015 Bitcoin spam campaign swelled the transaction pool to nearly 1 GB, crashing 10% of nodes, mostly memory-constrained like Raspberry Pi. Neonpool can easily withstand such an attack by generating additional bloom filters recursively as per demand, as described in section III.
@@
Replay transactions are rejected. Already seen transactions will trigger a positive in Neonpool indicating that the transaction is already present and thus will be dropped.
IV-F Summary
@@
The unconfirmed transaction pool plays a critical role in verifying, storing, and disseminating transactions while they await inclusion in a block. We present Neonpool, a novel transaction pool construction for cryptocurrencies that instead of storing complete transactions via map data structures, stores transaction fingerprints via bloom filters. We perform benchmarks using unique Bitcoin and Ethereum datasets comprising approximately 10 million unique transactions. We achieve up to two orders of magnitude reduction in memory consumption, fingerprinting up to 400 megabytes of data in as low as 2 MB, while maintaining a verification and forwarding accuracy exceeding 99.99%, with a slight increase in computation load. We also demonstrate its adversarial resilience. We summarize our findings in Tab. IV.
@@
Data Structure map-based mempool/txpool bloomtxFilter, dstxFilter
@@
TABLE IV: Neonpool vs Bitcoin/Ethereum
V Prior work
Researchers have proposed optimizations and light clients that reduce the computation, storage, and communication requirements of cryptocurrency nodes while ensuring robust security. Here we present prominent contributions in the literature, along with a comparison with Neonpool in Table V.
@@ V Prior work
EdraX[35] state Any Any New/Mod SparseMT, Dist.VC
@@
TABLE V: Comparison with other light-clients
Reducing blockchain overheads: Satoshi Nakamoto introduced Simplified Payment Verification (SPV) clients as a lightweight client, which requires download of only block headers and select blocks to verify transactions [10]. However, these scale linearly: Ethereum’s SPV client storage exceeds 10 GB as of July, 2023 [28].
@@
Reducing bootstrapping costs: Kiayias et al. [12] introduced sublinear storage complexity in SPV clients via skip lists, termed noninteractive proofs of proof-of-work (NIPoPoW). This solution checks for high-difficulty previous blocks. Verifying a logarithmic number of these suffices to ensure security for the whole chain. However, this solution is only practical in an honest network with fixed difficulty, unlike most cryptocurrencies that have variable block difficulty.
FlyClient [2] achieves logarithmic complexity, using Merkle Mountain Range Commitments to achieve memory improvements, and a random block sampling protocol to ensure security. This solution works even if parts of the network are adversarial and also with variable block difficulty. However, NiPoPoW and Flyclient still require linear resources, and verifying transactions remains costly, as for each verified transaction, the corresponding block must also be downloaded.
TXCHAIN [26] addresses this issue, using contingent transaction aggregation as a technique for compressing transaction inclusion proofs. Proof of Necessary Work [4] performs necessary computation for system verification within the proof-of-work computation, utilizing SNARKs and Pederson hash.
@@
Bloom filters in cryptocurrencies: Bloom filters have been employed for tasks such as generating block summaries and event logs, filtering transactions for SPV nodes, enabling efficient pool synchronization and block declaration, ensuring fairness in block selection, ascertaining unique addresses, and examining privacy concerns. We refer the reader to [6] for a summary of existing applications and some potential future applications of probabilistic data structures in blockchains.
Optimizing the transaction pool: This remains a neglected area in the research literature. In Bitcoin, the mempool is allocated 300 MB by default [38]. In Ethereum the default number of executable and non-executable transaction slots per account and globally are 16, 64, 4096, and 1024 respectively [19]. Surplus transactions are evicted. Users can define a custom transaction pool acceptance policy. In a low-memory environment, the transaction pool size can be disabled entirely.
Our work is a new direction in the domain of light clients. Being orthogonal to existing light clients, these may be integrated with Neonpool to maximize resource saving.
VI Future Work and Conclusion
@@
Our approach, to utilizing bloom filters, in principle, can be extended to other cryptocurrencies. It is straightforward to adapt Neonpool-BTC to UTXO-based and Neonpool-ETH to account-based cryptocurrencies, tweaking parameters such as expiry time and filter size according to network conditions.
Moving forward, we aim to conduct individual studies for cryptocurrencies, collecting data, documenting their architecture in-depth, and further exploring the trade-offs between error rates, memory utilization, reprocessing costs, security, and computation. We intend to initiate a Bitcoin Improvement Proposal (BIP) and an Ethereum Improvement Proposal (EIP).

### v2

@@ 4.4 Computation time
Query 1.8 1.6 2.9 4.3 2.1 1.9 3.4 5.1
Insert 6.9 1.8 3.3 5.1 8.2 2.1 3.8 5.8
Table 3: Query/Insert time ($\mu$s)
4.5 Security analysis
Here, we establish the security of Neonpool, focusing on two main aspects: 1. whether errors made by Neonpool compromise its security or that of the broader network; and 2. Neonpool’s resilience to adversarial attacks.
The network-level impact of false positives vanishes at the network level, i.e., effectively negligible. Each node initializes its Bloom filters with a unique, random 128-bit salt, ensuring independence between filters across nodes. Consequently, false positives at one node are statistically independent of those at other nodes. For instance, with a Bloom filter accuracy of 99.99% (corresponding to a false positive rate of 0.0001 per Neonpool node), the probability of two nodes erroneously dropping the same transaction is $(0.0001)^{2}$, an exceedingly low likelihood.
Furthermore, research indicates that while transaction pools across nodes are not entirely identical, they exhibit a remarkable 99% similarity in their contents [14]. This high consistency underscores that, despite occasional discrepancies introduced by Bloom filter false positives or false negatives, the overall integrity of transaction pool contents across the network remains robust, ensuring reliable transaction propagation.
In Bitcoin and Ethereum, there is an established practice of limiting transaction pool sizes and managing overflow by rejecting or expiring excess transactions. For instance, Bitcoin’s default mempool size is capped at 300 MB, while Ethereum employs a default limit of 4096 transactions in the transaction pool, with surplus transactions being evicted [47] [24]. Users can customize transaction pool policies or disable transaction pools entirely to accommodate low-memory or low-computation environments.
@@ 4.5 Security analysis
Additionally, an adversary may: 1. trigger false positives to censor specific transactions; 2. craft invalid transactions that evade verification and validation; 3. generate spam.
@@
Literature shows that any bloom filter can be efficiently transformed to be adversarial resilient by applying a pseudo-random permutation of the input [13] i.e. applying a sufficiently large (128-bit) random salt before forwarding it to the bloom filter. This change requires little overhead and randomizes the adversary’s queries by applying a pseudorandom permutation to them; then, we may consider the transactions sent by the attacker as random and not as chosen adaptively by the adversary. It is also recommended that a node regenerate its 128-bit random salt every time a new bloomtxFilter or dstxFilter is generated. Thus, the adversary only has oracle access to the bloom filter and does not know its contents or seed.
We situate our assumption within established practices in the cryptocurrency ecosystem and light client security models: Secure random seed generation is essential for the security of Bitcoin, Ethereum, and the broader crypto ecosystem. It underpins wallet key generation, ensuring unique private keys. Random seeds also facilitate contract deployment through ECDSA nonce generation, enable ephemeral session keys for secure communication, and support multi-signature wallets by creating unique key shares for participants.
As identified by Chatzigiannis et al. [18], there are several common assumptions that underpin light client designs, including trusted genesis block, reliable consensus, secure underlying cryptographic primitives, weak synchrony (i.e., no long network partitions), trusted setup, peer-to-peer communication for relaying information, and rational behaviour of participants. Secure underlying cryptographic primitives and trusted setups are of particular interest to us.
In this context, our assumption about Oracle access to the Bloom filter aligns with these principles. Secure generation and protection of the 128-bit random seed are essential for the Bloom filter’s adversarial resilience and are consistent with best practices in decentralized systems. The cryptographic strength and manipulation resistance of the hash functions in Bloom filters fundamentally depend on this secure initialization and randomization.
@@
Secondly, Eve may craft invalid transactions to evade verification and validation, i.e., attempt double-spending. Neonpool preserves the verification and validation mechanisms of Bitcoin and Ethereum. Regarding transactions with conflicting inputs or out-of-order nonce, typically, nodes accept and forward the first seen transaction, and the first seen can differ for nodes. It is the job of miners not to add conflicting transactions to a block and the network nodes to screen incoming blocks. Since Neonpool maintains complete UTXO and Trie information, Neonpool nodes will reject blocks that include double-spend transactions.
@@
Thirdly, Eve might launch a dust or spam attack or replay transactions. A 2015 Bitcoin spam campaign swelled the transaction pool to nearly 1 GB, crashing 10% of nodes, mostly memory-constrained like Raspberry Pi. Neonpool can withstand such attacks by recursively generating additional bloom filters on demand, as described in section 3.
@@
Replay transactions are rejected. Already seen transactions will trigger a positive in Neonpool, indicating that the transaction is already present and thus will be dropped.
4.6 Summary
@@
The unconfirmed transaction pool plays a critical role in verifying, storing, and disseminating transactions while they await inclusion in a block. We present Neonpool, a novel transaction pool construction for cryptocurrencies that stores transaction fingerprints via bloom filters instead of storing complete transactions via map data structures. We perform benchmarks using unique Bitcoin and Ethereum datasets comprising approximately 10 million unique transactions. We achieve up to two orders of magnitude reduction in memory consumption, fingerprinting up to 400 megabytes of data in as low as 2 MB while maintaining a verification and forwarding accuracy exceeding 99.99%, with a slight increase in computation load. We also demonstrate its adversarial resilience. We summarize our findings in Tab. 4.
Due to their function in the network, Neonpool node operators are not required to store full transactions. This conceptually resembles Bitcoin Core’s built-in "pruned node" option, which reduces hard disk requirements by storing only a few recent blocks instead of the full blockchain on disk while still contributing to the network’s footprint and health by validating and forwarding transactions.
Such full-node users operate nodes primarily to contribute to the Bitcoin network out of a sense of community or altruism, similar to how people operate nodes for the Tor network. Neonpool lowers the barrier to entry for such non-mining full nodes, offering them greater control over the memory resources they allocate for the transaction pool.
Neonpool provides full-node operators with significant flexibility to optimize resource usage by adjusting memory and computation allocations to balance efficiency and functionality. Lightweight devices, like IoT nodes, can store only transaction fingerprints in Bloom filters, while full nodes participating in mining can still retain a subset of full transactions to propose a block.
@@
Data Structure map-based bloomtxFilter,
mempool/txpool dstxFilter
@@
Table 4: Neonpool vs Bitcoin/Ethereum
5 Prior work
Our work relates to two main bodies of research: lightweight clients and transaction pool management.
5.1 Lightweight clients
Our work relates to two main bodies of research: lightweight clients and transaction pool management. Our work has a tangential relationship with existing light clients. While current light clients effectively address specific resource constraints, such as storage, computation, or bandwidth, none focus on reducing memory consumption in the transaction pool—a critical yet often overlooked challenge. Consequently, direct comparisons with these approaches are not feasible. However, this distinction helps position our work within the broader spectrum of light clients. Notably, these light-client solutions are orthogonal to our approach and can be deployed alongside Neonpool if needed. We present key contributions from the literature on lightweight clients, emphasizing that no existing solutions propose a lightweight version of the transaction pool.
@@ 5.1 Lightweight clients
EdraX[44] state Any Any New/Mod SparseMT,
Dist.VC
@@
Table 5: Neonpool in the light client spectrum
Reducing blockchain overheads: Satoshi Nakamoto introduced Simplified Payment Verification (SPV) clients as a lightweight client, which requires download of only block headers and select blocks to verify transactions [9]. However, these scales linearly: Ethereum’s SPV client storage exceeds 10 GB as of July 2023 [36].
@@
Reducing bootstrapping costs: Kiayias et al. [11] introduced sublinear storage complexity in SPV clients via skip lists, termed noninteractive proofs of proof-of-work (NIPoPoW). This solution checks for high-difficulty previous blocks. Verifying a logarithmic number of these suffices to ensure security for the whole chain. However, this solution is only practical in an honest network with fixed difficulty, unlike most cryptocurrencies with variable block difficulty.
FlyClient [2] achieves logarithmic complexity, using Merkle Mountain Range Commitments for memory improvements and a random block sampling protocol to ensure security. This solution works even if parts of the network are adversarial and have variable block difficulty. However, NiPoPoW and FlyClient still require linear resources, and verifying transactions remains costly, as each verified transaction also requires downloading the corresponding block.
TXCHAIN [34] addresses this issue using contingent transaction aggregation to compress transaction inclusion proofs. Proof of Necessary Work [3] performs necessary system verification within the proof-of-work computation, utilizing SNARKs and Pederson hash.
@@
5.2 Transaction pool
This remains a neglected area in the research literature, with earlier works focusing on mitigating spam, dust, and DDoS attacks by filtering malicious transactions.
Baqer et al. were the first to emphasize the transaction pool’s significance in their analysis of a 2015 Bitcoin stress test [7]. This attack expanded the pool to nearly 1 GB, reportedly causing 10% of Bitcoin nodes to crash. They classified 23% of transactions as spam using clustering techniques and proposed spam filtering. However, they warned about the risks of misclassifying legitimate transactions: even a 1-2% false positive rate could create a self-inflicted DoS attack. Additionally, attackers could manipulate transaction attributes to bypass filters if mechanisms were exposed.
Subsequent works focused on transaction filtering. Saad et al. introduced Contra, which filters spam based on age and fee thresholds [20]. Configuring these thresholds presents tradeoffs: high thresholds risk false positives, while low thresholds risk false negatives. Their modelling estimates 60% accuracy, recommending dynamically increasing block size to accommodate dust transactions and deter spammers, though results lack real-world validation. Wang et al. proposed an Anti-dust solution, analyzing Bitcoin transactions (2009-2017) with a Gaussian model [21]. Transactions below a threshold were classified as spam and placed in a dust pool, later moved to the mempool if space allowed. Overflowing transactions were discarded. Simulations showed dust transactions increased validation time from 200 to 25,000 seconds, while Anti-dust reduced it to 215 seconds.
Eduardo et al. simulated dust attacks on Ethereum using 2 million genuine transactions and synthetic ones [19]. They found dust attacks extended transaction pending time by over 42%. Using machine learning, they achieved 94% accuracy in identifying under-priced potential DoS attack transactions. DETER highlighted Ethereum-specific vulnerabilities, describing DETER-X and DETER-Z attacks that evict legitimate transactions, delay processing, and reduce miner revenue [22]. Proposed heuristics mitigate these attacks by regulating txpool entry and eviction.
To the best of our knowledge, Neonpool is the first optimization technique that re-architects the transaction pool of a cryptocurrency from an optimization perspective, specifically aiming to reduce the local memory consumption of the transaction pool. Earlier works primarily focus on mitigating spam, dust, and DDoS attacks by filtering out malicious or low-value transactions. They do this by introducing additional modules to the transaction pool while preserving the transaction pool structure itself. These approaches increase resource usage, as implementing real-time filtering (via statistical techniques or machine learning) or maintaining separate pools for spam imposes significant computational and memory costs, which are not adequately investigated.
Our work does not focus on spam filtering but prioritizes reducing memory usage and enhancing transaction pool resilience for larger traffic flows. As a result, a direct comparison with these techniques is not possible. Filtering solutions are orthogonal to our overall approach and can be deployed against Neonpool if required.
Our work is directly compared with the reference implementations of Bitcoin Core and Ethereum, as detailed in §4.
6 Conclusions and Future Work
@@
Our results highlight Neonpool’s potential and provide a foundation for further exploration in several directions.
Our work pioneers investigating the suitability of probabilistic data structures for transaction pool construction. While over a dozen Bloom filter variants exist in the literature, we begin with the simplest ones that meet our requirements and are widely understood. Future research will evaluate alternative probabilistic data structures, such as counting Bloom filters and cuckoo filters, to explore improved trade-offs in error rates, memory and computational overhead.
Another focus of future work is to assess Neonpool’s scalability under varying network conditions, transaction loads, and scenarios such as spam and dust attacks. Future experiments will aim to log and analyze attack patterns, examine the relationship between transaction fee spikes and network congestion, and design robust defences.
Finally, we plan to introduce Neonpool’s design to the cryptocurrency community by initiating a Bitcoin Improvement Proposal (BIP) and an Ethereum Improvement Proposal (EIP), to facilitate Neonpool’s live deployment and integration into existing blockchain ecosystems.
Neonpool’s approach can be extended to other cryptocurrencies with minimal modifications. For instance, Neonpool-BTC can be adapted for UTXO-based systems, and Neonpool-ETH for account-based cryptocurrencies by adjusting parameters like transaction expiry time and filter size based on network conditions and block time. These adaptations enable generic client-side optimizations, making Neonpool broadly applicable to lightweight devices across different distributed ledger technologies (DLTs). We are currently preparing datasets for alternative currencies, including Litecoin and Solana, to evaluate Neonpool’s adaptability and effectiveness across diverse blockchain architectures.